Operations Report April 25th

Crew 193 Operations Report 25Apr2018

SOL: 3

Name of person filing report: David Attig

Non-nominal systems: Com radio #6, Sim suits #10, 2, 5, 4, 6.

Notes on non-nominal systems: Radio #06 (newer style) only lasted 1 hour on EVA with a freshly charged battery. Radio will not be used for future EVAs. Sim suit 10 still has the badly scratched visor, suits 2, 5, 6 need maintenance on straps attaching neck ring to pack, and suit 4 has a broken charge connector.

Generator (time run): 2000-0715

Solar— SOC 95 % (Before generator is run at night)

Diesel – 75 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 0 gallons.

Water (trailer) – 3500 gallons.

Water (storage) – 550 gallons

Water (static) – 350 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134078.1

Toilet tank emptied: no
Date last emptied: April 24

ATV’s (Honda, 300, 350.1, 350.2, 350.3):
Honda Used?: Yes
300 Used? No
350.1 Used? Yes
350.2 Used? No
350.3 Used? No

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 1/4 Gallons

# Hours the ATVs were Used today: 1.2

Notes on ATVs: None

Deimos rover used: No
Hours: 124.0
Beginning charge: 100%
Ending charge: 100 %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 33.7
Beginning charge: 100 %
Ending charge: 100 %
Currently charging: No

Opportunity rover used: Out Of Service
Hours: 31
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Curiosity rover used: No
Hours: 23
Beginning charge: 100 %
Ending charge: 100 %
Currently charging: No

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: We would like the sim suit supply box dropped off in our airlock, so we can work on the two-piece suits.

Summary of internet: Appears to be working.

Summary of suits and radios: Radio 06 appears to transmit a beep continuously, draining the battery. Suits have some issues; see email titled “Details about EVA Suits”.

Summary of Hab operations: supported two EVAs and investigated suit status. Sol Summary report for more details.

Summary of GreenHab operations: plants are doing well, see GreenHab report for more details.

Summary of ScienceDome operations: Progress was made on dishwasher experiment, see Science Report for more details.

Summary of RAM operations: Placed tools in RAM for future use, took boards apart and painted some to use for an MDRS sign.

Summary of health and safety issues: Need to work on better balance between working in RAM and not breaking sim; see email “A few notes – about EVA duration” for details.

Questions, concerns and requests to Mission Support: A few requests, listed in Sol Summary.